# MoMentS: A Comprehensive Multimodal Benchmark for Theory of Mind

### v1

@@ Abstract
Understanding Theory of Mind is essential for building socially intelligent multimodal agents capable of perceiving and interpreting human behavior. We introduce MoMentS (Multimodal Mental States), a comprehensive benchmark designed to assess the ToM capabilities of multimodal large language models (LLMs) through realistic, narrative-rich scenarios presented in short films. MoMentS includes over 2,344 multiple-choice questions spanning seven distinct ToM categories. The benchmark features long video context windows and realistic social interactions that provide deeper insight into characters’ mental states. While the visual modality generally enhances model performance, current systems still struggle to integrate it effectively, underscoring the need for further research into AI’s multimodal understanding of human behavior.
@@
[…]
@@ 1 Introduction
To support the development of socially intelligent multimodal agents and assess current models’ ToM in realistic, socially grounded scenarios, we introduce MoMentS (Multimodal Mental States), a comprehensive multimodal video question-answering benchmark designed to evaluate ToM across seven abilities derived from the ATOMS taxonomy Beaudoin et al. (2020): Intentions, Desires, Beliefs, Knowledge, Percepts, Non-literal Communication, and Emotions. The dataset comprises 2,344 human-annotated questions and 9,376 candidate answers sourced from 168 long-form videos, annotated with short and long context windows, multimodal cue markers, and adversarially-generated distractors to minimize biases.
@@ 2 Related Work
Prior benchmarks for ToM broadly fall into two categories: text-only and multimodal. Traditional text-only benchmarks, such as TOMI Le et al. (2019) and HI-TOM He et al. (2023), predominantly focus on probing models’ ability for nested belief tracking and logical inference through text stories lacking realistic social context. TOMBench Chen et al. (2024) expands beyond belief tracking alone, incorporating a broader taxonomy of social and pragmatic ToM tasks (e.g., faux-pas detection, persuasion, hidden emotions, desires) within everyday textual scenarios. Despite this richer coverage, it remains constrained by its purely textual format, lacking multimodal information critical to human social understanding Byom and Mutlu (2013).
@@
Social Genome Mathur et al. (2025) (based on SocialIQ2 Wilf et al. (2023)) addresses the evaluation of social interaction understanding in VLMs through video-based multiple-choice questions, but videos are limited to 60 second clips, and evaluation is not designed to evaluate ToM. Furthermore, Guo et al. (2023) observed a strong bias in the representations of correct and incorrect answer candidates, where LLMs can achieve high accuracy with no context or transcription.
@@ 3.1 Taxonomy and Question Design
We adopt the ATOMS taxonomy (Abilities in Theory of Mind Space) introduced by Beaudoin et al. (2020) from their meta-analysis of ToM studies and proposed as a systematic framework for model evaluation by Ma et al. (2023). ATOMS categorizes ToM into seven distinct abilities: Knowledge, Emotions, Desires, Beliefs, Intentions, and Non-literal Communication (NLC). We describe each ability in Table 1. This taxonomy supports precise question formulation and provides a detailed framework for systematically evaluating specific ToM abilities in models. We design annotation guidelines (See Appendix A.9) around it.
@@
Knowledge Understanding what a person knows or does not know based on their sensory access. Why is the old man more interested in the box on the table than the book? The box is unfamiliar to him, therefore more interesting.
Emotions Identifying and reasoning about emotional responses, their evolution, and when emotions are hidden or complex. How is the girl feeling after hearing the news? She is saddened but hides her true feelings with a smile.
Desires Situations that involve preferences, conflicting desires, or actions driven by desire. Why is the man making an effort to defend his friend’s actions? He knows his friend stole, but feels attracted to him and wants to stay on good terms.
Beliefs Comprehending true and false beliefs and how beliefs influence actions. Why does the mother pretend to be surprised by the birthday cake? She thinks her daughter believes the cake is a surprise, even though she helped plan it.
Intentions Understanding goals, motivations, and the underlying reasons for actions. Why does the woman interrupt the conversation? She wants to change the topic.
Percepts Reasoning about what a character can or cannot perceive through their senses. Why does the old man not try to catch the falling vase? He did not see it, so he did not try to catch it.
NLC Interpreting humour, sarcasm, deception, and other speech that goes beyond literal meaning. Why does the man say he is stuck in traffic? He is lying to avoid conflict with his boss.
Table 1: Overview for ATOMS abilities covered in MoMentS with example question/answer pairs.
@@ 3.2 Video Selection
Existing datasets contain synthetic videos or minute-long clips to incorporate limited temporal context. We instead propose to use short films as these contain more complex characterizations and longer temporal contexts, while having a self-contained narrative. Our videos come from the SF20K dataset Ghermi et al. (2024), which contains a curated collection of short films from the YouTube channel Omeleto. Ghermi et al. (2024) verified that these films exhibit minimal information leakage to state-of-the-art language models compared to other common video sources like the sitcom Friends. Additionally, the videos are high-quality, vary in length (10 to 20 minutes), and provide complete, cohesive stories.
Not all short films have scenarios suitable for evaluating ToM. To filter these out, we prompted GPT-4o with film synopses to identify ones that likely contain interesting social dynamics. We then select videos with the highest likelihood of generating meaningful question-answer pairs and assign each annotator a a subset of these to annotate.
@@ 3.3 Data Organization
includes one correct option and three distractors. Annotators are instructed to write distractors that are as plausible as possible, such that only a nuanced understanding of the context can reveal the correct answer. We paid special attention to the distractors, see Section 4.2 for more details on this.
@@
  • Full Context Window $[t_{0},t_{j}]$: A longer segment starting from the beginning of the video, intended to provide full narrative context useful for understanding character backgrounds, motivations, and evolving social dynamics.
  • Focused Context Window $[t_{i},t_{j}]$: A shorter segment containing only the immediate context required to answer the question. This window excludes broader narrative information, focusing instead on the specific scene being queried.
During evaluation, we explicitly instruct models that the question refers to the end of the provided interval ($t_{j}$). This approach minimizes reliance on temporal references that may hint at the correct answer, which requires understanding the interaction. If leveraged effectively, the Full Context Window provides all the information required to understand characters, providing better insights into their mental states and interpersonal dynamics.
@@ 4 Annotation Methodology
(a) Boxplot of accuracies in the Focused Context Window.
(b) Boxplot of accuracies in the Full Context Window.
Figure 3: Boxplots comparing average accuracies across different models (at $\sim$7B parameter scale) and ToM abilities. Results for Audiovisual LLMs (▲) are reported using video and audio inputs ($VA$), while results for Video LLMs (⚫) use video and transcript inputs ($VT$).
@@ 4.1 Annotation Pipeline
$[t_{i},t_{j}]$ length $42.44\pm 55.5$
$[t_{0},t_{j}]$ length $388.47\pm 262.3$
Number of Videos $168$
@@
Table 2: Top: Mean length $\pm$ SD of questions, correct answers, and distractors (in words), together with the average duration of the focused $[t_{i},t_{j}]$ and full $[t_{0},t_{j}]$ context windows (in seconds). Bottom: Number of Videos and average duration (in minutes)
@@
The main annotation phase spanned six weeks and involved 16 annotators who collectively produced 2,344 questions. This phase followed the methodology developed during the second pilot and incorporated several design choices aimed at improving quality and reducing bias (see Figure 2 for an overview):
@@ 5 Experimental Evaluations
Model $[t_{0},t_{j}]$ $[t_{i},t_{j}]$
$T$ $VT$ $T$ $VT$
@@
$VA$
Human 86
Table 3: Accuracy of different models on MoMentS, reported for both Full $[t_{0},t_{j}]$ and Focused $[t_{i},t_{j}]$ context windows. For Video LLMs (⚫), we show performance with transcripts only (T) and video+transcripts (VT). For Audiovisual LLMs (▲), we additionally report audio (A) and video+audio (VA) inputs. For Speech LLMs (▼), we report performance using audio only (A). *Human evaluation was conducted on a subset of 100 samples in the Full $[t_{0},t_{j}]$ setting using video and audio (VA) by a participant external to the annotation process.
5.1 Experimental Setup
@@
These models process visual, text, and audio inputs. We evaluate Qwen2.5-Omni 7B Xu et al. (2025), VideoLLaMa2 7B Cheng et al. (2024), and MiniCPM-o 2.6 (8B) Yao et al. (2024).
@@
All models are evaluated using the transformers library Wolf et al. (2020) with temperature set to 0. Under 70B paramenters models run on a single NVIDIA A100 GPU, whereas larger models (>70B) run on three NVIDIA H100 GPUs.
@@ 5.2 LLM Evaluation
We evaluate models under different input conditions: textual or audio dialogues only (Transcripts ($T$) or Audio ($A$)), and combined inputs (Video+Transcripts ($VT$) or Video+Audio ($VA$)). Additionally, we compare model performance across two temporal contexts: the Full Context Window $[t_{0},t_{j}]$ and the Focused Context Window $[t_{i},t_{j}]$.
@@
Figure 3 presents boxplots with per-model scatter points, showing model accuracy across different ToM abilities under two context window conditions. As observed in the globlal accuracies, models perform better using shorter Focused Context Windows, though the impact of context length varies by ability. For example, accuracy is notably higher for Emotions and Intentions questions with shorter contexts, suggesting that these tasks rely more on immediate cues. Among results within the longer Full Context Window, Knowledge, Desires, and Non-literal Communication (NLC) questions perform relatively better, suggesting that longer context may be beneficial for understanding characters’ background and effectively answering these questions. For both of the context settings, Percepts and Beliefs remain the most challenging abilities. Future work should investigate how context window length affects human performance in this task. In Figure 4, we further analyze the role of visual input by averaging accuracies across Video LLMs. While the visual modality is not yet fully leveraged, it still contributes positively across all abilities, indicating its importance for ToM reasoning.
Figure 4: Average accuracies across Video LLMs highlighting the effect of the vision modality accuracy on different ToM abilities (Focused Context Window).
@@
$\Delta_{t_{j}-t_{0}}$ $\Delta_{VT-T}$ $\Delta_{VA-A}$
Body Langauge 3.71 7.92 8.05
@@
Table 4: Effect of context length and visual input on questions marked as reliant on multimodal cues. $\Delta_{t_{j}-t_{0}}$ is the average accuracy difference between Focused and Full Contexts, with positive values indicating better performance in shorter intervals. $\Delta_{VT-T}$ and $\Delta_{VA-A}$ represent average accuracy gains from adding visual input (VT vs. T across Video and Audiovisual LLMs; VA vs. A across Audiovisual LLMs), both reported for the shorter context interval.
We further analyze performance on questions requiring multimodal understanding (facial expression or gaze, body language, and speech-related cues). As shown in Table 4, incorporating visual input and using a shorter context window generally improves performance, particularly for questions involving Body Language and Facial Expressions or Gaze. Speech-related questions benefit less from visual information when dialogues are provided as transcripts ($VT$), indicating stronger reliance on textual cues; however, vision seems to show a larger benefit in these questions when dialogues are provided directly as audio inputs in audiovisual models ($VA$).
@@ 5.3 Evaluation on Answer Set Bias
In this section, we evaluate the impact of using an LLM-in-the-loop design during the annotation pipeline, specifically for distractor creation. For MCQA-style ToM evaluation to be meaningful, questions should not be answerable without access to some form of context such as video, audio, or transcripts. However, as observed in our initial pilot and in prior work Guo et al. (2023), models often exploit biases in question-answer sets to guess the correct answer even without contextual input.
@@
By reducing biases in the answer sets, we create greater headroom for models to improve through reasoning rather than shortcut learning.
@@ 6 Open Challenges for Future Model Development
Answering Who said what, when? requires time-synchronised links between each utterance, the speaking character, and the surrounding visual context. Without such alignment, models cannot track which speakers possess which knowledge, nor can they exploit gaze, facial expressions, or body language that modulate dialogue meaning. The small gains we observe from adding vision (Table 3), and the limited improvements on questions marked as reliant on visual cues (Table 4), indicate that existing pipelines underutilize this channel.
@@
Uniform frame sampling risks missing short yet meaningful signals while wasting computation on redundant content. Simply increasing the frame rate is expensive and, as our ablation in Appendix A.5 shows, does not improve performance. Specialized frame sampling strategies that prioritise human-salient events (faces, hands, gaze shifts) are needed to capture the cues that observers actually rely on.

### v2

@@ Abstract
Understanding Theory of Mind is essential for building socially intelligent multimodal agents capable of perceiving and interpreting human behavior. We introduce MoMentS (Multimodal Mental States), a comprehensive benchmark designed to assess the ToM capabilities of multimodal large language models (LLMs) through realistic, narrative-rich scenarios presented in short films. MoMentS includes over 2,300 multiple-choice questions spanning seven distinct ToM categories. The benchmark features long video context windows and realistic social interactions that provide deeper insight into characters’ mental states. We evaluate several MLLMs and find that although vision generally improves performance, models still struggle to integrate it effectively. For audio, models that process dialogues as audio do not consistently outperform transcript-based inputs. Our findings highlight the need to improve multimodal integration and point to open challenges that must be addressed to advance AI’s social understanding.
@@
Emilio Villa-Cueva1, S M Masrur Ahmed2, Rendi Chevi1, Jan Christian Blaise Cruz1, Kareem Elzeky1, Fermin Cristobal1, Alham Fikri Aji1, Skyler Wang3, Rada Mihalcea4, Thamar Solorio1 1MBZUAI, 2University of Houston, 3McGill University, 4University of Michigan github.com/villacu/MoMentS
@@ 1 Introduction
To support the development of socially intelligent multimodal agents and assess current models’ ToM in realistic, socially grounded scenarios, we introduce MoMentS (Multimodal Mental States), a comprehensive multimodal video question-answering benchmark designed to evaluate ToM across seven abilities derived from the ATOMS taxonomy Beaudoin et al. (2020): Intentions, Desires, Beliefs, Knowledge, Percepts, Non-literal Communication, and Emotions. The dataset comprises 2,335 human-annotated questions and 9,340 candidate answers sourced from 168 long-form videos, annotated with short and long context windows, multimodal cue markers, and adversarially-generated distractors to minimize biases.
@@ 2 Related Work
Prior benchmarks for ToM broadly fall into two categories: text-only and multimodal. Traditional text-only benchmarks, such as TOMI Le et al. (2019) and HI-TOM Wu et al. (2023), predominantly focus on probing models’ ability for nested belief tracking and logical inference through text stories lacking realistic social context. TOMBench Chen et al. (2024) expands beyond belief tracking alone, incorporating a broader taxonomy of social and pragmatic ToM tasks (e.g., faux-pas detection, persuasion, hidden emotions, desires) within everyday textual scenarios. Despite this richer coverage, it remains constrained by its purely textual format, lacking multimodal information critical to human social understanding Byom and Mutlu (2013).
@@
Social Genome Mathur et al. (2025) (based on SocialIQ2 Wilf et al. (2023)) addresses the evaluation of social interaction understanding in VLMs through video-based multiple-choice questions, but videos are limited to 60 second clips, and evaluation is not designed to evaluate ToM. Furthermore, Guo et al. (2023) observed a strong bias in the representations of correct and incorrect answer candidates, where LLMs can achieve high accuracy with no context at all.
@@ 3.1 Taxonomy and Question Design
We adopt the ATOMS taxonomy (Abilities in Theory of Mind Space) introduced by Beaudoin et al. (2020) from their meta-analysis of ToM studies and proposed as a systematic framework for model evaluation by Ma et al. (2023). ATOMS categorizes ToM into seven distinct abilities: Knowledge, Emotions, Desires, Beliefs, Intentions, and Non-literal Communication (NLC). We describe and exemplify each ability in Table 1. This taxonomy supports precise question formulation and provides a detailed framework for systematically evaluating specific ToM abilities in models. We design annotation guidelines (See Appendix A.10) around it.
@@
Knowledge Understanding what a person knows or does not know based on their sensory access. Why is the soldier interested in the boy’s bottle? The soldier does not know what is inside the bottle and wants to find out.
Emotions Identifying and reasoning about emotional responses, their evolution, and when emotions are hidden or complex. How do the old woman and the young woman feel in this conversation? The younger woman feels annoyed, and the older woman feels angry.
Desires Situations that involve preferences, conflicting desires, or actions driven by desire. What does the girl want after walking past the group and reading the sign? She wants to go in to the establishment, finding it appealing.
Beliefs Comprehending true and false beliefs and how beliefs influence actions. What does the woman with the ponytail think of the man who is watching TV? She thinks the man watching TV is aggressive.
Intentions Understanding goals, motivations, and the underlying reasons for actions. Why does the old man give a beer to the bearded man and leave the cabin? He wants the bearded man to follow him so they can talk outside.
Percepts Reasoning about what a character can or cannot perceive through their senses. Why didn’t the woman with the long hair protect herself from the man? Because the man came up behind her and she didn’t see him.
NLC Interpreting humour, sarcasm, deception, and other speech that goes beyond literal meaning. Why does the young man in white ask the man in blue if he likes his work? He’s being sarcastic and wants to annoy the man in blue.
Table 1: Overview for ATOMS abilities with example question/answer pairs extracted from MoMentS
@@ 3.2 Video Selection
Existing datasets contain synthetic videos or minute-long clips that provide short temporal context. We instead propose to use short films as these contain more complex characterizations and longer temporal contexts, while having a self-contained narrative. Our videos come from the SF20K dataset Ghermi et al. (2025), which contains a curated collection of short films from the YouTube channel Omeleto. Ghermi et al. (2025) verified that these films exhibit minimal information leakage to state-of-the-art language models compared to other common video sources like the sitcom Friends. Additionally, the videos are high-quality, vary in length (10 to 20 minutes), and provide complete, cohesive stories.
Not all short films have scenarios suitable for evaluating ToM. To filter these out, we prompted GPT-4o with film synopses to identify ones that likely contain interesting social dynamics. We then select videos with the highest likelihood of generating meaningful question-answer pairs and assign each annotator a subset of these to annotate.
@@ 3.3 Data Organization
includes one correct option and three distractors. Annotators are instructed to write distractors that are as plausible as possible, such that only a nuanced understanding of the context can lead to the correct answer. We paid special attention to the distractors, see Section 4.2 for more details on this.
@@
  • Full Context Window $t_{0},t_{j}]$: A longer segment starting from the beginning of the video, intended to provide full narrative context useful for understanding character backgrounds, motivations, and evolving social dynamics.
  • Focused Context Window $t_{i},t_{j}]$: A shorter segment containing only the immediate context required to answer the question. This window excludes broader narrative information, focusing instead on the specific scene being queried.
During evaluation, we explicitly instruct models that the question refers to the end of the provided interval ($t_{j}$). This approach minimizes reliance on temporal references in the question that may hint at the correct answer, which requires understanding the interaction. If leveraged effectively, the Full Context Window provides all the information required to understand characters, providing better insights into their mental states and interpersonal dynamics.
@@ 4.1 Annotation Pipeline
$t_{i},t_{j}]$ length (s) $42.44\pm 55.5$
$t_{0},t_{j}]$ length (s) $388.47\pm 262.3$
Number of Videos 168168
@@
Table 2: Top: Mean length $\pm$ SD of questions, correct answers, and distractors (in words), together with the average duration of the Focused $t_{i},t_{j}]$ and Full $t_{0},t_{j}]$ Context Windows (in seconds). Bottom: Number of Videos and average duration (in minutes)
@@
The main annotation phase spanned six weeks and involved 16 annotators who collectively produced 2,335 questions. This phase followed the methodology developed during the second pilot and incorporated several design choices aimed at improving quality and reducing bias (see Figure 2 for an overview):
@@ 5 Experimental Evaluations
5.1 Experimental Setup
Model $t_{0},t_{j}] [$$t_{i},t_{j}]$
$T$ VTVT $T$ VTVT
@@ 5.1 Experimental Setup
VAVA
*Human (average of 3) 86.33 $\pm$ 1.15
Human (majority-vote) 91.0
Table 3: Accuracy of different models on MoMentS, reported for both Full $t_{0},t_{j}]$ and Focused $t_{i},t_{j}]$ context windows. For Video LLMs (⚫), we show performance with transcripts only (T) and video+transcripts (VT). For Audiovisual LLMs (▲), we additionally report audio (A) and video+audio (VA) inputs. For Speech LLMs (▼), we report performance using audio only (A). *Human evaluation is carried out in a sample of 100 randomly drawn items.
@@
These models process visual, text, and audio inputs. We evaluate Qwen2.5-Omni 7B Xu et al. (2025), VideoLLaMa2 7B Cheng et al. (2024), and MiniCPM-o 2.6 (8B) Yao et al. (2024). We observed that MiniCPM-o 2.6 yielded errors in long-form videos; therefore, for this model we only report results on the Focused Context Window.
@@
All models are evaluated using the transformers library Wolf et al. (2020) with temperature set to 0. Under 70B parameters models run on a single NVIDIA A100 GPU, whereas larger models (>70B) run on three NVIDIA H100 GPUs.
@@ 5.2 LLM Evaluation
We evaluate models under different input conditions: textual or audio dialogues only (Transcripts ($T$) or Audio ($A$)), and combined inputs (Vision+Transcripts ($VT$) or Vision+Audio ($VA$)). Additionally, we compare model performance across two temporal contexts: the Full Context Window $t_{0},t_{j}]$ and the Focused Context Window $t_{i},t_{j}]$.
@@
As a reference, we conducted human evaluations on a subset of 100 randomly sampled questions with three different evaluators on the Full Context Window and the $VA$ setting. Individual accuracies were 87.0/85.0/87.0 (an average of 86.3). Majority-vote accuracy with ties marked as incorrect was 91.0. We observed a percent agreement of 0.80, and a Fleiss $\kappa$ of 0.733, which indicates substantial agreement between evaluators when selecting an answer. In Table 3, we report both majority-vote (marking ties as incorrect) and the average of individual accuracies.
(a) Boxplot of accuracies in the Focused Context Window.
(b) Boxplot of accuracies in the Full Context Window.
Figure 3: Boxplots comparing accuracies across different models (at $\sim$7B parameter scale) and ToM abilities. Results for Audiovisual LLMs (▲) are reported using video and audio inputs (VAVA), while results for Video LLMs (⚫) use video and transcript inputs (VTVT).
@@
Figure 3 presents boxplots with per-model scatter points, showing model accuracy across different ToM abilities under two context window conditions. As observed in the global accuracies, models perform better using shorter Focused Context Windows, though the impact of context length varies by ability. For example, accuracy is notably higher for Emotions and Intentions questions with shorter contexts, suggesting that these tasks rely more on immediate cues. Among results within the longer Full Context Window, Knowledge, Desires, and Non-literal Communication (NLC) questions perform relatively better, suggesting that longer context may be beneficial for understanding characters’ background and effectively answering these questions. For both of the context settings, Percepts and Beliefs remain the most challenging abilities. Future work should investigate how context window length affects human performance in this task.
In Figure 4, we compare the effect of visual input for Video and Audiovisual LLMs. Vision contributes positively across all abilities, although the improvement varies by model type. For Video LLMs, the gains are relatively consistent, while Audiovisual LLMs show greater variation, especially when dialogues are provided as audio: some abilities improve by over 12 points, others by only 6-8. This larger gain with vision is mainly due to the weaker performance of the audio-only setting, where models struggle on abilities such as Beliefs or Intentions that likely require more complex information that the visual channel can provide. Notably, Non-literal Communication shows the smallest improvement, suggesting stronger reliance on dialogue than visual cues compared to the other abilities.
(a) Video LLMs ($\sim$7B scale)
(b) Audiovisual LLMs (VTVT vs. $T$)
(c) Audiovisual LLMs (VAVA vs. AA)
Figure 4: Comparison between average accuracies across the evaluated Video LLMs and Audiovisual LLMs with and without vision across different ToM abilities (Focused Context Window). The number in parentheses refers to the improvement due to the visual modality.
@@
$\Delta_{Focus-Full}$ $\Delta_{VT-T}$ $\Delta_{VA-A}$
Body Language 3.71 7.92 8.05
@@
Table 4: Effect of context length and visual input on questions marked as reliant on multimodal cues. $\Delta_{Focus-Full}$ is the average accuracy difference between Focused and Full Contexts, with positive values indicating better performance in shorter intervals. $\Delta_{VT-T}$ and $\Delta_{VA-A}$ represent average accuracy gains from adding visual input (VT vs. T across Video and Audiovisual LLMs; VA vs. A across Audiovisual LLMs), both reported for the shorter context interval.
We further analyze performance on questions requiring multimodal understanding (facial expression or gaze, body language, and speech-related cues). As shown in Table 4, incorporating visual input and using a shorter context window generally improves performance, particularly for questions involving Body Language and Facial Expressions or Gaze. Speech-related questions benefit less from visual input when dialogues are provided as transcripts ($VT$), confirming stronger reliance on text. Interestingly, Audiovisual models show larger gains in Speech-related questions when dialogues are provided as audio ($VA$), suggesting vision is more beneficial in this setting.
@@ 5.3 Evaluation on Answer Set Bias
In this section, we evaluate the effectiveness of using an LLM-in-the-loop design during the annotation pipeline, specifically for distractor creation. For MCQA-style ToM evaluation to be meaningful, questions should not be answerable without access to some form of context such as video, audio, or transcripts. However, as observed in our initial pilot and in prior work Guo et al. (2023), models often exploit biases in question-answer sets to guess the correct answer even without contextual input.
@@
By reducing biases in the answer sets, we create greater headroom for models to improve through reasoning based on the provided context.
@@ 6 Open Challenges for Future Model Development
Prior work has shown that state-of-the-art models struggle to attribute utterances to speakers in multimodal conversations Chang et al. (2025). Answering Who said what, when? requires time-synchronized links between each utterance, the speaking character, and the surrounding visual context. Without such alignment, models cannot track which speakers possess which knowledge, nor can they exploit gaze, facial expressions, or body language that modulate dialogue meaning. The small gains we observe from adding vision (Table 3), and the limited improvements on questions marked as reliant on visual cues (Table 4), also indicate that existing pipelines underutilize this channel.
@@
Uniform frame sampling risks missing short yet meaningful signals while wasting computation on redundant content. Simply increasing the frame rate is expensive and, as our ablation in Appendix A.5 shows, does not improve performance. Specialized frame sampling strategies that prioritize human-salient events (faces, hands, gaze shifts) are needed to capture the cues that observers actually rely on.